# An empirical appraisal of methods for the dynamic prediction of survival with numerous longitudinal predictors

### v1

@@ 3 Datasets and implementation
Sections 3.1-3.3 present the three datasets considered in this article. For each dataset we describe how the data were gathered, the length of follow-up and the frequency of repeated measurements; then, we define the survival outcome that we want to predict, and list the baseline and longitudinal covariates that are employed to predict it. As it will become apparent later in this section, the three datasets differ substantially in sample size, length of the follow-up, frequency and number of repeated measurements per subject and number of longitudinal covariates employed as predictors, making them representative of different scenarios that one may encounter when dealing with real-world longitudinal studies.
Lastly, in Section 3.4 we provide information about the implementation of the dynamic prediction methods for these datasets, including information about landmark times and prediction horizons, performance measures, tuning parameters and further computational details.
@@ 3.1 The ADNI dataset
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) study (Weiner et al.,, 2010) is an ongoing study started in 2004 that was designed to identify and validate biomarkers related to the progression of AD. By April 2023, the study had enrolled 2428 individuals, each of which is scheduled to undergo an initial visit and subsequent assessments scheduled at 3, 6, 12, 18, 24, 36, and 48 months from baseline, followed by annual visits thereafter. Throughout the follow-up visits participants undergo comprehensive evaluations for dementia, including a variety of cognitive assessments, as well as biospecimen sampling and brain imaging analysis.
Differently from the ROSMAP study that will be introduced in Section 3.2, where the cause of dementia (AD or not AD) is reported, the ADNI study only reports dementia diagnoses, without attributing them to AD or other causes. At each visit, ADNI participants are classified into one of the following categories: cognitive normal (CN), mildly cognitive impaired (MCI), or affected by dementia. Moreover, both baseline information about the participant and longitudinal measurements of cognitive, imaging and biochemical markers are collected in this study.
Our analysis aims to predict the time from baseline until a dementia diagnosis for subjects that entered the study as CN or MCI. To predict this survival outcome we employ 5 baseline covariates (age, gender, baseline diagnosis, number of apolipoprotein $\epsilon$4 alleles, and number of years of education) and 21 longitudinal covariates that are listed in Supplementary Table 1. These covariates were selected based on their relevance as predictors of dementia and on the fact that they did not have too many missing values.
After removing subjects already diagnosed with dementia at baseline, subjects without any follow-up information after baseline, and subjects with missing covariate values at baseline the baseline visit, 1643 subjects were retained for analysis. The number of visits per subject varies from 1 to 22, with a mean of 6.2 visits. The follow-up period ranges between 3 months and 15.5 years, with a mean of 3.4 years. Supplementary Figure 1 illustrates how number of individuals at risk changes over time, and the Kaplan-Meier estimator of the probability to be free from dementia in the ADNI dataset.
@@ 3.2 The ROSMAP dataset
Every year, subjects enrolled in ROSMAP undergo a clinical assessment that leads to an evaluation of their cognitive status, which is classified into 6 classes: no cognitive impairment (NCI); mild cognitive impairment (MCI); MCI and another condition contributing to cognitive impairment (MCI+); Alzheimer’s dementia (AD); AD and other condition contributing to CI (AD+); other primary cause of dementia, without clinical evidence of AD (Other). In addition, longitudinal information about a wide range of genomic, experiential, psychological and medical risk factors is collected alongside information (such as age, gender, education) about the participant’s background.
In this article we focus on predicting the time (measured from the date of study enrollment) until AD is diagnosed (i.e. , cognitive status AD or AD+) for subjects that entered the study in the NCI, MCI, MCI+ and Other categories. To do so we employ 5 baseline covariates (age, gender, years of education received, baseline cognitive status and presence of a cancer diagnosis) and a set of 30 longitudinal covariates, listed in Supplementary Table 2, which measure aspects related to several different domains (these covariates were selected from a larger list of longitudinal covariates both because of their relevance as risk factors for AD, and because they did not have too many missing values).
After removing subjects already diagnosed with AD at baseline, subjects without any follow-up information after baseline, and subjects with missing covariate values at baseline, 3293 subjects were retained for analysis. The number of visits per subject varies from 2 to 30, with a mean of 9.3 visits. The follow-up period ranges between 1 and 29 years, with a mean of 8.3 years. Supplementary Figure 2 illustrates how number of individuals at risk evolves over time, and the Kaplan-Meier estimator of the probability to be free from AD in the ROSMAP dataset.
@@ 3.3 The PBC2 dataset
The PBC2 dataset contains data from a clinical trial on primary biliary cholangitis (PBC) conducted between 1974 and 1984 at the Mayo Clinic (Murtaugh et al.,, 1994). The randomized trial involved 312 participants who were randomized between a treatment group where patients received the drug D-penicillanime and a placebo group. The trial recorded the occurrence of the first of two survival outcomes: liver transplantation, or death. Because most of the methods considered in our comparison cannot deal with competing risks, for the purpose of our analysis we focus on predicting time to death, treating patients who underwent liver transplantation as right-censored at the date of transplantation.
To predict time to death, we employ 3 baseline covariates (age, gender and treatment group) alongside with the 8 longitudinal covariates listed in Supplementary Table 3. Patients enrolled in the trial underwent visits upon study entry, 6 and 12 months after baseline, and yearly visits after that. The number of follow-up visits per patient ranged between 1 and 22, with an average of 6.1; the follow-up period ranged from 0.1 to 14.1 years, with a mean of 4.6 years. Supplementary Figure 3 shows the number of patients at risk and the Kaplan-Meier estimator of the survival probability for the PBC2 dataset.
3.4 Application of the dynamic prediction methods to the ROSMAP, ADNI and PBC2 datasets
@@ 3.4.1 Definition of the landmark and horizon times
The datasets considered in our study differ substantially in their sample size, frequency of the repeated measurements (visits) and length of the follow-up. All of these aspects affect both the number of individuals still at risk at a given landmark, and the number of repeated measurements per subject available up to the landmark time. Lastly, the length of the follow-up across patients affects the possibility to evaluate predictive performance at a given horizon.
For these reasons, we determined the landmark and horizon times at which predictions of survival are evaluated separately for each dataset, trying to balance the need to have a good number of repeated measurements to model the evolution of the longitudinal covariates with MFPCA or LMMs, and that of having a good number of individuals still at risk to be able to estimate the Cox / RSF model and reliably assess predictive performance. Clearly, this is easier for the ADNI and ROSMAP datasets, whose sample sizes are considerably bigger than that of PBC2.
@@ 3.4.2 Performance measures and validation of predictive performance
To evaluate the predictive accuracy of the different models we computed the following performance measures for survival data: the Brier score (Graf et al.,, 1999), the time-dependent AUC (Heagerty et al.,, 2000), and the C index (Pencina and D’Agostino,, 2004). We employed repeated cross-validation (CV) to obtain unbiased estimates of the predictive performance of the different models. For the ROSMAP and ADNI datasets we employed 10-fold CV with 10 repetitions; for the PBC2 dataset, where using 10-fold CV led to very small validation folds that would have resulted in unreliable estimates of predictive performance, we employed 5-fold CV with 20 repetitions.
@@ 3.4.3 Implementation details
In order to guarantee a fair comparison of the methods included in the analysis, all methods were trained on data from the same set of subjects. Specifically, we excluded from the datasets subjects for which either (a) the value of one or more baseline covariates was missing or (b) the value of one of the longitudinal predictors at the baseline visit was missing to avoid the need of multiple imputation that, given the small number of subjects for which this happened, would have unnecessarily complicated our analyses. While (a) is important for all methods, (b) was added as filtering criterion to ensure that there are no missing values when estimating the static Cox model.
In addition, the 4 methods that model the evolution of the longitudinal predictors over time required additional preprocessing.
Both MFPCCox and FunSRF use MFPCA to reduce the longitudinal covariates to MFPCA scores. The estimation of such scores relies on the so-called ”PACE algorithm”, which estimates the univariate FPCA scores from each longitudinal variable by assuming that the measurements have been obtained using a grid that is regular across subjects. Unfortunately, in the real-world datasets that we consider such assumption is not met, as often visits do not happen exactly at the pre-planned time but may take place a few days / months earlier or later. To overcome this issue, we “regularized” the measurement times across subjects by aligning all visits to the preplanned visit schedule of the specific dataset.
DynForest and PRC employ LMMs to summarize the longitudinal covariates into predicted random effects. Estimation of LMMs can often lead to convergence problems when the response variable is strongly skewed. This can be problematic in our method comparison, where we need to be able to estimate these LMMs several times during the repeated CV. To reduce the frequency of such convergence problems and the need to address convergence issues with ad-hoc solutions each time they arise, we proceeded to apply a log-transformation to those longitudinal covariates whose skewness index was larger than 1, and a cubic transformation to those with skewness index below -1.
Furthermore, many of the methods involved in our comparison require the specification of one or more tuning parameters:
@@
  • PRC allows to choose between the ridge, elasticnet and lasso penalty. We employed the ridge penalty (pencal’s default), and used the pre-implemented cross-validation to determine the value of the tuning parameter of the ridge penalty;
  • the RSF within FunRSF requires the specification of several parameters. The random forests were trained using an ensemble of $B=1000$ trees. The hyperparameters of each tree were selected taking into account the specificities of each dataset. The minimum number of subject in a terminal node $s$ was required to be at least 15 for all the datasets. The number of predictors $F$ to be randomly selected at each node was set to the square root of the total number of predictors available in the given dataset;
  • the random forests within DynForest were trained using $B=200$ trees as estimating DynForest is much more time-consuming than estimating FunSRF. Setting the tuning parameters $F$, $s$ and $e$ for this method is more complicated than for FunRSF, as the need to re-estimate the LMMs after each node split often leads to convergence problems. For this reason, multiple values of $s$ were sequentially tried in case of convergence problems (Supplementary Table 4), progressively increasing $s$ to reduce the chance of encountering convergence errors towards the end of the tree. The value of $F$ was set in the same way as with FunRSF; $e$ was set equal to 5 for ROSMAP and ADNI, and to 4 for PBC2.
@@ 4 Results
4.1 ADNI
The repeated CV estimates of the C index, tdAUC and Brier score for the dynamic prediction of time to dementia in the ADNI study are presented in Table 1 and in Figures 1 and 2.
C index. As shown in Table 1, for all landmark times PRC achieves the highest performance, and FunRSF the lowest one. The relative performance of the other methods slightly changing across landmarks: at landmarks 2 and 3, landmarking and static Cox have the second and third highest C index, and are followed by DynForest and MFPCCox. Instead, at landmark 4 DynForest is the second best method. Overall, the relative performance of DynForest and MFPCCox improves as the landmark time increases, whereas that of landmarking and static Cox worsens.
@@ 4.1 ADNI
Table 1: Cross-validated C index estimates (and standard deviation) for the ADNI dataset
@@
Time-dependent AUC. Overall, the results for the tdAUC (Figure 1) are very much in line with the ones for the C index: PRC is again the best performing method, and FunRSF the worse. Landmarking is the second best at landmark 2, and DynForest at 4; the two methods have similar performance at landmark 3. MFPCCox and FunRSF typically show lower accuracy; compared to the other methods, the performance of LOCF landmarking and of the Static Cox model tends to worsen as the landmark increases.
@@
Brier score. In general, the difference in performance across methods appears to be smaller when considering the Brier score (Figure 2). At all landmarks, the static Cox model performs quite well for predictions until year 7; from year 8, its performance starts to worsen in comparison to other methods. PRC is usually the best or second best performing method, followed by landmarking, DynForest and MFPCCox. Once again, FunRSF performs worse than all other methods.
@@
  1. A1)
    landmarking and the static Cox model show overall a decent performance, specially at earlier landmark times. Surprisingly, these methods often outperform MFPCCox, and always outperform FunRSF;
  2. A2)
    methods that rely on mixed-effects models, namely PRC and DynForest, typically outperform the MFPCA-based methods (MFPCCox and FunRSF);
  3. A3)
    conditionally on the method (LMM, or MFPCA) used to summarize the longitudinal predictors, the methods that use a Cox model for the survival outcome (PRC and MFPCCox) outperform the corresponding methods that use a RSF (DynForest and FunRSF);
  4. A4)
    overall, the approach used to model the longitudinal predictors has a stronger effect on predictive performance than the approach used to model the survival outcome.
4.2 ROSMAP
The repeated CV estimates of the C index, tdAUC and Brier score for the application to the ROSMAP data are presented in Table 2 and in Figures 3 and 4. At landmarks 2 and 5, no results were produced for DynForest due to the fact that estimation of this model failed on all or most (100 and 97 / 100, respectively) folds during the repeated CV. Similarly, at landmark 4 no results are presented for MFPCCox and FunRSF as computation of the MFPCA scores failed on 97 of the 100 folds. These convergence issues are due to a higher percentage of missing visit / values in this dataset, that complicates the modelling of the longitudinal predictors.
C index. PRC and LOCF landmarking appear to be the best performing methods according to the estimates of the C index (Table 2). The difference in performance between the two methods is negligible at landmark 2, and slightly higher at later landmarks. MFPCCox ranks third at landmarks 2 to 4, and fourth at landmark 6; DynForest ranks fourth at landmarks 3 and 4, and third at landmark 6. FunRSF and the Static Cox model exhibit the worse predictive performance.
@@ 4.2 ROSMAP
Table 2: Cross-validated C index estimates (and standard deviation) for the ROSMAP dataset.
@@
Time-dependent AUC. The tdAUC estimates show that the difference in performance across methods is larger at the later landmark times. At landmark 2, PRC and landmarking achieve the highest tdAUC, followed by MFPCCox. At landmark 3, DynForest has the highest tdAUC at prediction times 4 to 6, and PRC at times 7 to 15. At landmark 4, PRC has the highest tdAUC at all prediction times and is followed by DynForest, landmarking and MFPCCox. At landmark 6, PRC, landmarking and DynForest clearly outperform the other 3 methods.
@@
Brier score. Similarly to the tdAUC, also for the Brier score estimates (Table 4) the difference in performance across methods increases with the landmark. At all landmarks, PRC, landmarking,MFPCCox and DynForest outperform the static Cox model and FunRSF. The difference between these methods is more visible at landmarks 3 and 4, where PRC has the lowest Brier score, DynForest the highest, and landmarking and MFPCCox lay in between the two.
Summary. Differently from the ADNI dataset, in this case the Static Cox model exhibits a much worse predictive performance. This seems to indicate that for the ROSMAP dataset, using repeated measurement data can improve predictive performance considerably. Once again, landmarking exhibits a good predictive performance, usually ranking between second and fourth best performing method. PRC is typically first, and MFPCCox and DynForest between second and fourth depending on the performance measure and landmark.
4.3 PBC2
The repeated CV estimates of the C index, tdAUC and Brier score for the PBC2 dataset are presented in Table 3 and in Figures 5 and 6.
C index. At all landmark times, PRC achieves the highest value of the C index (Table 3); surprisingly, at landmarks 2.5 and 3 the static Cox model is the second best performing method according to this metric, and at landmark 3.5 the third. Landmarking and DynForest are next, followed by MFPCCox and, last, FunRSF. The standard error of the C index estimates for this dataset is typically larger than for the ADNI and ROSMAP datasets: a result, this, that can be ascribed to the smaller sample size available in the PBC2 study to evaluate predictive performance.

### v2

@@ 3 Datasets and implementation
Sections 3.1-3.3 present the three datasets used for the benchmarking. We describe how the data were gathered, the length of follow-up and the frequency of repeated measurements, define the survival outcome that we want to predict, and list the baseline and longitudinal covariates that are employed to predict it. The datasets differ substantially in sample size, length of the follow-up, frequency and number of repeated measurements per subject, and number of longitudinal covariates employed as predictors (Supplementary Table 1), making them representative of different scenarios that one may encounter when dealing with real-world longitudinal studies.
Lastly, in Section 3.4 we provide information about the implementation of the dynamic prediction methods for these datasets, including information about landmark times and prediction horizons, performance measures, and modelling and computational details.
@@ 3.1 The ADNI dataset
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) study (Weiner et al., , 2010) is an ongoing study started in 2004 that was designed to identify and validate biomarkers related to the progression of AD. By April 2023, the study had enrolled 2428 individuals, each scheduled to undergo an initial visit and subsequent assessments scheduled at 3, 6, 12, 18 and 24 months from baseline, followed by annual visits thereafter. At each follow-up visit, participants undergo comprehensive evaluations for dementia, including a variety of cognitive assessments, biospecimen sampling and brain imaging analysis.
Differently from the ROSMAP study that will be introduced in Section 3.2, where the cause of dementia (AD or not AD) is reported, the ADNI study only reports dementia diagnoses, without attributing them to AD or other causes. At each visit, ADNI participants are classified into one of the following categories: cognitive normal (CN), mildly cognitive impaired (MCI), or affected by dementia. Moreover, both baseline information about the participant and longitudinal measurements of cognitive, imaging and biochemical markers are collected.
Our goal is to predict the time until a dementia diagnosis for subjects that entered the study as CN or MCI. To predict this survival outcome we employ 5 baseline covariates (age, gender, baseline diagnosis, number of apolipoprotein $\epsilon$4 alleles, and number of years of education) and 21 longitudinal covariates that are listed in Supplementary Table 2 (these covariates were selected based on their relevance as predictors of dementia, provided that they did not have too many missing values). After removing subjects already diagnosed with dementia at baseline, subjects without any follow-up information after baseline, and subjects with missing covariate values at the baseline visit, 1643 subjects were retained for analysis. The number of visits per subject varies from 1 to 22, with a mean of 6.2 visits. The follow-up length ranges from 3 months to 15.5 years, with a mean of 3.4 years. Supplementary Figure 1 illustrates how the number of individuals at risk evolves over time, and the Kaplan-Meier estimator of the probability to be free from dementia in this dataset.
@@ 3.2 The ROSMAP dataset
Every year, subjects enrolled in ROSMAP undergo a clinical assessment that leads to an evaluation of their cognitive status, which is classified into 6 classes: no cognitive impairment (NCI); mild cognitive impairment (MCI); MCI and another condition contributing to cognitive impairment (MCI+); Alzheimer’s dementia (AD); AD and other condition contributing to CI (AD+); other primary cause of dementia, without clinical evidence of AD (Other). In addition, longitudinal information about a wide range of genomic, experiential, psychological and medical risk factors is collected alongside information about the participant’s background (such as age, gender and education).
We aim to predict the time until AD is diagnosed (i.e., cognitive status AD or AD+) for subjects that entered the study in the NCI, MCI, MCI+ and Other categories. To do so, we employ 5 baseline covariates (age, gender, years of education received, baseline cognitive status and presence of a cancer diagnosis) and a set of 30 longitudinal covariates, listed in Supplementary Table 3, which measure aspects related to several different domains (these covariates were selected from a wider list of longitudinal covariates based on their relevance as risk factors for AD, provided that they did not have too many missing values). After removing subjects already diagnosed with AD at baseline, subjects without any follow-up information after baseline, and subjects with missing covariate values at baseline, 3293 subjects were retained for analysis. The number of visits per subject varies from 2 to 30, with a mean of 9.3 visits. The follow-up period ranges from 1 to 29 years, with a mean of 8.3 years. Supplementary Figure 2 illustrates how the number of individuals at risk evolves over time, and the Kaplan-Meier estimator of the probability to be free of AD in the ROSMAP dataset.
@@ 3.3 The PBC2 dataset
The PBC2 dataset contains data from a clinical trial on primary biliary cholangitis (PBC) conducted between 1974 and 1984 at the Mayo Clinic (Murtaugh et al., , 1994). The randomized trial involved 312 participants who were randomized between a placebo group, and a treatment group where patients received the drug D-penicillanime. The trial recorded the occurrence of the first of two survival outcomes: liver transplantation, and death. Because most of the multi-step methods considered in our comparison cannot deal with competing risks, for the purpose of our analysis we focus on predicting time to death, treating patients who underwent liver transplantation as right-censored at the date of transplantation.
To predict time to death, we employ 3 baseline covariates (age, gender and treatment group) alongside the 8 longitudinal covariates listed in Supplementary Table 4. Patients enrolled in the trial underwent visits upon study entry, 6 and 12 months after baseline, and yearly visits after that. The number of follow-up visits per patient ranged from 1 to 22, with an average of 6.1; the follow-up period ranged from 0.1 to 14.1 years, with a mean of 4.6 years. Supplementary Figure 3 shows the number of patients at risk and the Kaplan-Meier estimator of the survival probability for the PBC2 dataset.
3.4 Application of the dynamic prediction methods to the ADNI, ROSMAP and PBC2 datasets
@@ 3.4.1 Definition of the landmark and horizon times
The datasets considered in our study differ substantially in their sample size, frequency of the repeated measurements (visits) and length of the follow-up. All these aspects have an influence on the number of individuals still at risk at a given landmark, and the number of repeated measurements per subject available up to the landmark time. Lastly, the length of the follow-up across patients affects the possibility to evaluate predictive performance at a given horizon time.
Given all these differences, we determined the landmark and horizon times at which predictions of survival are evaluated separately for each dataset, striving to balance the need to have a good number of repeated measurements to model the evolution of the longitudinal covariates with MFPCA or LMMs, and that of having a good number of individuals still at risk to be able to estimate the Cox / RSF model and reliably assess predictive performance. Clearly, this is easier for the ADNI and ROSMAP datasets, whose sample sizes are considerably bigger than that of PBC2.
@@ 3.4.2 Performance measures and validation of predictive performance
We employed the Brier score (Graf et al., , 1999), time-dependent AUC (Heagerty et al., , 2000), and C index (Pencina and D’Agostino, , 2004) as measures of predictive performance. We employed repeated cross-validation (RCV) to obtain unbiased estimates of the predictive performance of the different models. For the ROSMAP and ADNI datasets we employed 10-fold cross-validation (CV) with 10 repetitions; for the PBC2 dataset, where using 10-fold CV led to very small validation folds, we employed 5-fold CV with 20 repetitions.
@@ 3.4.3 Implementation details
To ensure that every method was trained on data from the same set of subjects, in each dataset we excluded subjects for which (a) the value of one or more baseline covariates was missing or (b) the value of one of the longitudinal predictors at the baseline visit was missing. Condition (a) is needed by all methods to ensure that there are no missing values in the baseline covariates, whereas (b) aims to ensure that are no missing values when estimating the static Cox model (where we use the measurements of the longitudinal covariates at the baseline visit as baseline covariates). Besides this data selection step, the four multi-step methods required some additional preprocessing and modelling choices that are specific to each method.
@@
DynForest and PRC employ LMMs to summarize the longitudinal covariates into predicted random effects. When the response variable is strongly skewed, this can often lead to estimation (convergence) problems. To reduce the frequency of such convergence problems and the need to address convergence issues with ad-hoc solutions each time they arise during the RCV, we proceeded to apply a log-transformation to those longitudinal covariates whose skewness index was larger than 1, and a cubic transformation to those with skewness index below -1.
PRC allows to choose between the ridge, elasticnet and lasso penalty for the estimation of the Cox model in equation (9). In this study we used the ridge penalty, which is the default option in the pencal package and the recommended choice in Signorelli et al., (2021).
Lastly, the RSF algorithms used by FunRSF and DynForest require the specification of multiple parameters. As concerns FunRSF, the random forests were trained using an ensemble of $B=1000$ trees. The minimum number of subject in a terminal node $s$ was required to be at least 15 for all datasets. The number of predictors $F$ to be randomly selected at each node was set to the square root of the total number of predictors available in the given dataset. The random forests within DynForest were trained using $B=200$ trees because estimating DynForest is much more time-consuming than estimating FunSRF (see Section 4.4). Choosing the $F$, $s$ and $e$ parameters for this method is more complicated than for FunRSF, because the need to re-estimate the LMMs after each node split often leads to convergence problems within DynForest. To reduce the frequency of convergence problems during the RCV, multiple values of $s$ (listed in Supplementary Table 5) were used sequentially in case of convergence problems, progressively increasing $s$ to reduce the chance of encountering convergence errors towards the end of the tree. The value of $F$ was set in the same way as for FunRSF; $e$ was set equal to 5 for ROSMAP and ADNI, and to 4 for PBC2.
@@ 4 Results
We now present the results of our benchmarking study. In Sections 4.1, 4.2 and 4.3 we compare the predictive performance of the multi-step dynamic prediction methods on the ADNI, ROSMAP and PBC2 datasets, respectively. In Section 4.4, we turn our attention to computing time.
4.1 Predictive performance on the ADNI data
The RCV estimates of the C index, tdAUC and Brier score for the dynamic prediction of time to dementia in the ADNI study are presented in Table 1 and Figures 1 and 2.
@@ 4.1 Predictive performance on the ADNI data
Table 1: Cross-validated C index estimates (and standard deviation) for the ADNI dataset. For each landmark, the top two methods are highlighted in bold.
@@
Time-dependent AUC (Figure 1). The results for the tdAUC align closely with the C index. PRC is again the best performing method, and FunRSF the worse, at each landmark time. Landmarking is the second best at landmark 2, and DynForest at 4; the two methods have similar performance at landmark 3. MFPCCox and FunRSF typically show lower accuracy. Once again we can observe that when compared to the multi-step methods, the performance of LOCF landmarking and of the Static Cox model tends to worsen as the landmark increases.
@@
Brier score (Figure 2). The differences across methods appear to be smaller when looking at the Brier score. At all landmarks, the static Cox model performs quite well for predictions until year 7; from year 8, its performance starts to worsen in comparison to other methods. PRC is usually the best or second best performing method, followed by landmarking, DynForest and MFPCCox. Once again, FunRSF performs worse than all other methods at every landmark time.
@@
  1. 1.
    landmarking and the static Cox model have a good predictive performance at earlier landmark times (especially at $\ell=2$). Surprisingly, these methods often outperform MFPCCox, and systematically outperform FunRSF;
  2. 2.
    the two multi-step methods that rely on mixed-effects models (PRC and DynForest) typically outperform those that use MFPCA (MFPCCox and FunRSF);
  3. 3.
    conditionally on the method (LMM, or MFPCA) used to summarize the longitudinal predictors, the methods that use a Cox model for the survival outcome (PRC and MFPCCox) outperform the corresponding methods that use an RSF (DynForest and FunRSF);
  4. 4.
    overall, the approach used to model the longitudinal predictors has a stronger impact on predictive performance than the approach used to model the survival outcome.
4.2 Predictive performance on the ROSMAP data
The RCV estimates of the C index, tdAUC and Brier score for the application to the ROSMAP data are presented in Table 2 and Figures 3 and 4. At landmarks 2 and 5, no results were produced for DynForest due to the fact that estimation of this model failed on all or most (100 and 97 / 100, respectively) folds during the repeated CV. Similarly, at landmark 5 no results are presented for MFPCCox and FunRSF as computation of the MFPCA scores failed on 97 of the 100 folds. These convergence issues are due to a higher percentage of missing visit / values in this dataset than in the ADNI dataset, which complicates the modelling of the longitudinal predictors.
@@ 4.2 Predictive performance on the ROSMAP data
Table 2: Cross-validated C index estimates (and standard deviation) for the ROSMAP dataset. For each landmark, the top two methods are highlighted in bold.
C index (Table 2). PRC and LOCF landmarking appear to be the best performing methods at all landmark times. The difference in performance between the two methods is negligible at landmark 2, and somewhat larger at all other landmarks. MFPCCox ranks third at landmarks 2, 3 and 4, and fourth at landmark 6; DynForest ranks fourth at landmarks 3 and 4, and third at landmark 6. FunRSF and the Static Cox model exhibit the worse predictive performance at all landmarks.
@@
Time-dependent AUC (Figure 3). When looking at the tdAUC estimates, we notice that the performance difference across methods is larger at the later landmark times. Differently from the ADNI data, here we can observe that all multi-step methods and landmarking improve over the static Cox model, suggesting that updating prediction based on the longitudinal data may be more beneficial for the ROSMAP data. We can observe that PRC, DynForest and LOCF landmarking are the top performing methods across all landmarks, with PRC usually outperforming the other two methods on the later horizon times. MFPCCox is usually the fourth best performing method, followed by FunRSF and the static Cox model.
@@
Brier score (Figure 4). Similarly to the tdAUC, also for the Brier score estimates the difference in performance across methods increases with the landmark. Moreover, the relative performance of the methods is fairly consistent across landmarks. PRC and LOCF landmarking exhibit the lowest Brier scores, followed by MFPCCox and DynForest. Once again, FunRSF and the static Cox model have the worst predictive performance
Summary. On the ROSMAP data, incorporating repeated measurements data in the RPM seems to improve predictive performance more than in the ADNI dataset. The relative performance of the prediction models is somewhat similar to the one observed with the ADNI data, with PRC and LOCF landmarking performing particularly well, followed by DynForest and MFPCCox. Once again, FunRSF appears to be the worst performing method among the multi-step methods considered in this benchmarking.
4.3 Predictive performance on the PBC2 data
The RCV estimates of the C index, tdAUC and Brier score for the PBC2 dataset are presented in Table 3 and Figures 5 and 6.